OpenStudy (anonymous):

A square has a vertex at (0, 4) and a perimeter between 10 and 25 units. Which of the following could not be the remaining coordinates of this quadrilateral? Thanks!

OpenStudy (anonymous):

0, -3), (7, -3) and (7, 4) (0, -1), (5, -1), and (5, 4) (0, 0), (4, 0), and (4, 4) (0, 1), (3, 1), and (3, 4)

OpenStudy (anonymous):

Take a better look at your coordinates. (-2, 1) and (-2, 6) (4, 6) and (4, 1) Notice that the X coordinate values are the same for these pairs. The sides that represent these coordinates are 5 units in length. Now take a look at them arranged like this. (4, 1) and (-2, 1) (4, 6) and (-2, 6) This time, they are the same Y coordinates. The distance between these are 6. Therefore, your perimeter is 2(5) + 2(6) = 10 + 12 = 22 units. Hope this helps!